# Local Mechanisms of Compositional Generalization in Conditional Diffusion

### v2

@@ 3.1 Equivalence between compositional structure and local scores
Lemma 1 (Local conditional score is exact for conditional projective composition).
Let $p$ be a pixel-space CPC (Definition 2) with disjoint sets $\{M_{j}\}$. Let $s^{t}$ be an LCS (Definition 1) with subsets:
$L_{i}^{t}(\mathcal{J})$ $=\begin{cases}\{j\}\cap\mathcal{J},\quad\text{if }i\in M_{j}\\ \emptyset,\quad\text{else},\end{cases}\text{ and}\quad N_{i}^{t}=\begin{cases}M_{j},\quad\text{if }i\in M_{j}\\ M_{b},\quad\text{else},\end{cases}\text{ where }M_{b}:=M_{\mathcal{J}_{\text{all}}}^{\complement}.$
Then $s^{t}$ is exactly the score of $p^{t}$:
@@ 3.2 Feature-space conditional projective composition
What if CPC does not hold in pixel-space – as is typically the case for non-location conditioners, like the color-conditioning of Figure 5, or text-to-image prompts such as “a watercolor of a cat eating sushi with chopsticks”? Pixel-space CPC is unlikely since each condition potentially affects many pixels (e.g., “watercolor” style would apply to all pixels). In these cases, we hypothesize that the local unconditional denoising mechanism still applies at low noise (Kamb & Ganguli, 2024; Niedoba et al., 2024). But is there any hope of compositional generalization at high noise?
@@
It follows directly from Lemma 1 that if a distribution has a CPC structure in feature-space then its score is an LCS in feature-space:
Corollary 1 (LCS is exact for CPC in feature-space; informal).
@@
Suppose that $p(x|c)$ is a CPC in an appropriate feature-space: that is, $\mathcal{A}\sharp p(z|c)$ is a CPC, where $\mathcal{A}$ is an invertible transform, and $z:=\mathcal{A}(x)$ is the feature-space representation. Then the feature-space score $\nabla_{z}\log(\mathcal{A}\sharp p)^{t}(z|c)$ is an LCS with neighborhoods $N_{i},L_{i}$ related to the CPC subsets $\{M_{j}\}$ of $\mathcal{A}\sharp p$ as in Lemma 1.
@@
For example, an appropriate feature-space mapping might disentangle the concepts “watercolor”, “cat”, and “sushi” into the required CPC/LCS structure – even though these concepts interact in pixel space. (Note that defining LCS in terms of “sparse dependencies” rather than literal “locality” is essential in feature-space, where conditioners are not locations and interacting features need not be contiguous). However, challenges remain: if the scores are LCS in feature-space, in order to exploit the sparse structure, the model must learn the associated feature-space mapping and its inverse, in addition to the local subsets, making the learning process significantly more challenging. [^2] In fact, even learning the feature-space transform and its inverse for the noiseless distribution is not enough – the model technically needs to learn a feature-map for every noise level. In practice it may be approximately sufficient to learn a single mapping and its inverse, though this is not entirely clear. This argument is made precisely in Appendix D. We will use the term disentanglement as a shorthand for CPC/LCS structure within a learned feature space in the rest of the paper.
As a practical heuristic for identifying disentanglement, we adapt a lemma of Bradley et al. (2025). In this context, it states that a necessary, but not sufficient, condition for CPC in a feature-space given by $\mathcal{A}$ is orthogonality of the mean difference vectors in feature-space, i.e.
$(\mu_{i}-\mu_{b})^{T}(\mu_{j}-\mu_{b}),\quad\forall i\neq j,\quad\text{where }\mu_{i}:=\mathbb{E}_{(\mathcal{A}\sharp p)(z|c_{i})}[z],\quad\mu_{b}:=\mathbb{E}_{(\mathcal{A}\sharp p)(z|\emptyset)}[z].$ (3)
To evaluate heuristic 3, we can construct a cosine similarity matrix $\{d_{i}d_{j}/\|d_{i}\|\|d_{j}\|\}_{i,j}$, with $d_{i}=\mu_{i}-\mu_{b}$: low similarity off-diagonal is evidence of feature-space disentanglement.
Figure 5: Length generalization in color-conditioned CLEVR. (Left) We test length generalization in a CLEVR model conditioned on colors (rather than locations), trained on 1-5 objects. We test up to 8 colors (blue, brown, cyan, gray, green, purple, red, yellow), and find generalization up to 7 colors, suggesting that compositional structure may exist in a learned feature-space. (Right) We visualize disentanglement between individual colors in the network’s internal representation via the cosine similarity heuristic 3 (low similarity off-diagonal indicates disentanglement). Colors seem to be significantly disentangled in early encoder layer activations – which may form a compositional “feature-space” potentially promoting generalization.
Figure 6: Preliminary SDXL pixel-space locality study. An SDXL generated image with the prompt “a beautiful photograph with a horse in the middle, a dog on the left, and a cat on the right.” Heatmaps show pixel gradient magnitude at locations marked with ++. The conditional gradient magnitude w.r.t. individual words is also evaluated at the indicated pixel, with dominant words (if any) shown in red. SDXL shows some degree of pixel-locality (particularly at low noise) and some degree of conditional-locality (more so at higher noise); consistent with apparently only some degree of compositional generalization (note the failure to draw a cat).
Figure 7: Preliminary evidence for feature-space compositionality in SDXL. (Top): Disentanglement between concepts (dog, horse, cat, van Gogh, Monet, Hokusai, sushi, croissant) via heuristic 3 cosine similarities. Pixel-space lacks meaningful structure, VAE-space exhibits some structure, and a clearer structure emerges within a proposed compositional Feature-space comprised of U-net mid block activations, showing higher intra-group similarity (e.g. dog, horse, cat) and lower inter-group similarity (e.g. cat, van Gogh). Compositions show that within multi-concept prompts (e.g. cat+sushi+van Gogh), individual token (e.g. cat) activations have high similarity with corresponding single-concept (cat) activations, consistent with compositional structure. (Middle): Disentanglement heuristic across all U-net cross-attention blocks (larger version in Figure 13). (Bottom): SDXL example generations. Disentangled concepts in the proposed feature-space (e.g. cat+sushi+van Gogh) compose more successfully than highly-entangled concepts (e.g. horse+dog+cat).
@@ 4 Additional Experiments
Our theory shows an equivalence between LCS and CPC (which implies length generalization). We test this directly in location-conditioned CLEVR models, where the compositional structure holds in pixel-space, and location-conditioners possess a direct notion of locality. Further, we test whether LCS could be a causal mechanism via a direct intervention: enforcing an explicitly LCS architecture to “fix” length generalization that previously failed. Turning to feature-space compositionality, we show partial length generalization in color-conditioned CLEVR and connect it with feature-space disentanglement. Finally, we make preliminary investigations of local/compositional structure in both pixel- and feature-space in SDXL.
@@
Figure 2 shows pixel- and conditional-locality in Experiments 1 and 2 (we will discuss Experiment 3 separately). We first observe that both models maintain pixel-locality at both low and high noise levels, contrasting with prior empirical findings on datasets like CIFAR-10 (Kamb & Ganguli, 2024; Niedoba et al., 2024) – reproduced in Figure 12 – showing delocalization at high noise. The difference likely stems from CLEVR images containing multiple, nearly-independent objects; unlike the datasets with a single centered subject studied in prior work. Second, we note significant differences in conditional-locality between Experiments 1 and 2. The length-generalizing Exp. 1 model exhibits strong conditional-locality at high noise, while the non-length-generalizing Exp. 2 model lacks conditional-locality at high noise (scores near conditioned locations either fail to respond to any conditions or depend on several non-local conditions). At low noise, both models transition to pixel-local unconditional denoising, as in Kamb & Ganguli (2024); Niedoba et al. (2024). These experiments support our prediction that length generalization depends on pixel- and conditional-locality, and suggest that conditional-locality at high noise plays a particularly important role. Details in Appendix F.
@@ A Causal Intervention in location-conditioned CLEVR
Experiment 3 (Figure 1) tests our theory via a direct causal intervention, wherein we design a local model architecture that explicitly enforces local conditional scores (as confirmed in Figure 2). We train the local model using the same conditioning as the failing Exp. 2 (labeling only a single object location), and find that the local architectural intervention restores length generalization (Table 1), “fixing” the failure. In fact, the local model trained on only one object can length-generalize up to 6 locations, unlike even the Exp. 1 model (Figure 4, Table 1). This supports the hypothesis that local conditional scores could be a causal mechanism for compositional generalization. Additional details in Section E.3.
@@
To better understand feature-space compositionality, we explore length generalization in color-conditioned CLEVR in Figure 5 and Table 1. For color-conditioning, we would expect compositional structure to exist in an appropriate feature space that disentangles the different colors. We observe that length generalization is possible to some extent – e.g. a model trained on 1-5 objects can generalize to 7 colors. Next, we investigate whether the model actually learns a disentangled feature space (i.e. one with CPC/LCS structure) by analyzing layer activations using the cosine similarity heuristic 3, and find evidence of color-disentanglement within several layers (Figure 5 right). Details in Section E.4.
A preliminary SDXL investigation
Can local mechanisms help to explain compositional generalization in real-world diffusion text-to-image models? We begin to explore this question by studying compositional/local structure in both pixel- and feature-space in a pretrained (‘out-of-the-box’) SDXL model (Podell et al., 2023) model. In Figure 6, we investigate pixel-space locality, by choosing a prompt that contains implicit location information (“middle”, “right”, “left”). We find some degree of locality even at high noise, and somewhat sparse dependencies on the text conditioner (measured by splitting the prompt into individual words and computing the score delta as we ablate each word). For example, the region where the horse ends up being drawn has the strongest dependency on the word “horse” at most noise levels, similar to observations in Chefer et al. (2023); Hertz et al. (2022). The local structure is far from perfect, but notably so is the compositional generalization: it fails to draw a cat as requested in the prompt. Details in Appendix G and F.
Next, we move beyond pixel-space to investigate whether SDXL may learn a disentangled feature-space. Certain concepts, such as animals vs. art styles, are almost certainly entangled in pixel-space but might be disentangled in the network’s internal representation. We analyze U-net cross-attention layer activations using the cosine similarity heuristic 3, and find evidence of CPC/LCS structure within several layers of the midblock: related concepts like dog, cat, and horse have higher similarity, while concepts like cat and van-Gogh have lower similarity. We also show examples connecting this feature-space structure with successful and unsuccessful compositions. Details in Appendix G.
@@ 5 Discussion and future work
Whether local mechanisms can explain compositional generalization in real-world diffusion models remains an open question. The preliminary SDXL experiments in Section 4 are meant only to be suggestive, and much more work is needed to fully understand compositional generalization in modern text-to-image models. This setting presents several challenges. First, since we often don’t know what was in the training set, it is unclear which prompts are actually OOD. Second, some kinds of compositional structure, such as style+content, exist only in feature-space, requiring more complex studies of the model’s learned representation as in Figures 5 and 7, potentially relying on heuristics such as 3. Despite significant evidence for feature disentanglement in diffusion representations (see Appendix A), this is still not fully understood. Future work could also exploit our finding of local conditional scores as a mechanism to improve compositional generalization. Our causal intervention in Experiment 3 shows that in a simple setting, enforcing an explicitly local architecture can improve generalization, suggesting that similar architectural, training, or inference-based interventions might be able to improve real-world models. Several existing methods can be viewed as implicit “local interventions”: for instance, layout-to-image methods that use explicit local constraints or biases (see Appendix A), or more generally, sparse attention architectures Child et al. (2019); Sun et al. (2022). Our theory helps to explain why these approaches are beneficial, and suggests more precise ways to target compositional generalization by specifically enforcing local conditional scores. When compositional structure exists only in feature-space, such interventions become more complex. The challenge becomes two-fold: we must first identify – or attempt to induce – feature-space transformations that reveal the compositional structure, and then apply local interventions within the learned feature space. Larger-scale studies on complex, real-world datasets are essential to clarify the challenges and explore opportunities to improve compositional generalization via local interventions.

### v3

@@ 3.1 Equivalence between compositional structure and local scores
Lemma 1 (Equivalence of CPC and LCS).
A distribution $p(x|c_{\mathcal{J}})$ is a pixel-space Conditional Projective Composition (Definition 2) with disjoint sets $\{M_{j}\}_{j\in\mathcal{J}_{\text{all}}}$ if and only if its score $s(x|c_{\mathcal{J}})$ is a Local Conditional Score (Definition 1) with subsets
$L_{i}^{t}(\mathcal{J})$ $=\begin{cases}\{j\}\cap\mathcal{J},\quad\text{if }i\in M_{j}\\ \emptyset,\quad\text{else},\end{cases}\text{ and}$
@@ 3.2 Feature-space conditional projective composition
What if CPC does not hold in pixel-space – as is typically the case for non-location conditioners, like text-to-image prompts such as “a watercolor of a cat eating sushi with chopsticks”? Pixel-space CPC is unlikely since each condition potentially affects many pixels (e.g., “watercolor” style would apply to all pixels). In these cases, we hypothesize that the local unconditional denoising mechanism still applies at low noise (Kamb and Ganguli, 2024; Niedoba et al., 2024). But is there any hope of compositional generalization at high noise?
@@
It follows directly from Lemma 1 that if a distribution has a CPC structure in feature-space then its score is an LCS in feature-space (we name these F-CPC/F-LCS, respectively):
Corollary 1 (F-LCS is exact for F-CPC; informal).
@@
Suppose that $p(x|c)$ is an F-CPC (a CPC in feature-space): that is, $\mathcal{A}\sharp p(z|c)$ is a CPC, where $\mathcal{A}$ is an invertible transform, and $z:=\mathcal{A}(x)$ is the feature-space representation. Then the feature-space score $\nabla_{z}\log(\mathcal{A}\sharp p)^{t}(z|c)$ is an F-LCS (an LCS in feature-space) with neighborhoods $N_{i},L_{i}$ related to the CPC subsets $\{M_{j}\}$ of $\mathcal{A}\sharp p$ as in Lemma 1.
@@
For example, in an appropriate feature-space, the concepts “watercolor”, “cat”, and “sushi” might have F-CPC structure – despite interacting in pixel space. (Note that an F-LCS will usually have sparse-dependencies rather than literal “locality”, as allowed by Definition 1, since interacting features need not be contiguous). However, challenges remain: if the scores are F-LCS in some feature-space, in order to exploit the sparse structure the model must learn this feature-space mapping and its inverse, in addition to the local subsets, making the learning process significantly more challenging. [^2] In fact, even learning the feature-space transform and its inverse for the noiseless distribution is not enough: the model technically needs to learn a noise-dependent feature-map, since pixel-space noise does not commute with the feature-space transform. This argument is made precisely in Appendix D.
As a practical heuristic for identifying F-LCS structure, we propose an empirically-testable necessary-but-not-sufficient condition for F-LCS, based on orthogonality between score differences (similar to Bradley et al. (2025) Lemma 8.1). The proof is in Section D.1.
Lemma 2 (F-LCS necessary-but-not-sufficient heuristic).
Let $s_{\mathcal{A}}^{t}(z|c):=\nabla_{z}\log(\mathcal{A}\sharp p)^{t}(z|c)$ be an F-LCS score in a feature-space given by transform $\mathcal{A}$, and let $t_{\max}$ denote the highest noise level. Then:
$d_{i}^{T}d_{j}$ $=0\quad\forall i\neq j,\quad d_{i}:=\mathbb{E}[s_{\mathcal{A}}^{t_{\max}}(\cdot|c_{i})]-\mathbb{E}[s_{\mathcal{A}}^{t_{\max}}(\cdot|\emptyset)]$
where the expectation is over the feature-space transformed noise distribution $\mathcal{A}\sharp\mathcal{N}(0,\sigma_{t_{\max}})$.
Practically, to study a feature-space represented within layer $\ell$ of a denoising network, we compute $s_{\mathcal{A}}^{t_{\max}}(\cdot|c)$ by drawing a noise sample, running the first denoising step (at time $t_{\max}$) to compute the conditional score, and hooking the activation of layer $\ell$. To obtain $d$ we average over multiple noise samples and compute the conditional-unconditional difference. Finally, we can construct a cosine similarity matrix $\{d_{i}d_{j}/\|d_{i}\|\|d_{j}\|\}_{i,j}$: low similarity off-diagonal is evidence of F-LCS.
Remark 1.
F-CPC/F-LCS structure should be viewed as a type of disentanglement, on which there is a rich literature (Bengio et al., 2013; Higgins et al., 2017; Chen et al., 2018; Kim and Mnih, 2018; Locatello et al., 2019; Kotovenko et al., 2019; Locatello et al., 2019; Watters et al., 2019; Yang et al., 2023; Zhang et al., 2023). To quote Karras et al. (2019): “There are various definitions for disentanglement, but a common goal is a latent space that consists of linear subspaces, each of which controls one factor of variation.” F-CPC/F-LCS satisfies this definition with the additional requirement that the subspaces be orthogonal. Intuitively, disentanglement is often thought to promote compositionality; our specific definitions and theory of F-CPC/F-LCS makes this connection precise and provable.
Remark 2.
Identifying feature-space disentanglement is fundamentally difficult since independence between high-dimensional random variables cannot be tested in polynomial time. However, a variety of practical metrics have been proposed (Higgins et al., 2017; Kim and Mnih, 2018; Chen et al., 2018); Locatello et al. (2019) shows that many common metrics are fairly correlated with each other. The heuristic of Lemma 2 is part of the broader family of closely-related disentanglement metrics, but is specifically designed to test F-CPC/F-LCS.
@@ 4 Additional Experiments
Our theory shows an equivalence between LCS and CPC (which implies length generalization). We test this directly in location-conditioned CLEVR models, where the compositional structure holds in pixel-space, and location-conditioners possess a direct notion of locality. Further, we test whether LCS could be a causal mechanism via a direct intervention: enforcing an explicitly LCS architecture to “fix” length generalization that previously failed. We also show partial length generalization in color-conditioned CLEVR. Finally, we investigate local/compositional structure in both pixel- and feature-space in SDXL, providing quantitative evidence for local conditional scores in feature-space.
@@
Figure 2 shows pixel- and conditional-locality in Experiments 1, 2, and 3. We first observe that Exp. 1 and 2 maintain pixel-locality at both low and high noise levels, contrasting with prior empirical findings on datasets like CIFAR-10 (Kamb and Ganguli, 2024; Niedoba et al., 2024) – reproduced in Figure 11 – showing delocalization at high noise. The difference likely stems from CLEVR images containing multiple, nearly-independent objects; unlike the datasets with a single centered subject studied in prior work. The non-length-generalizing Exp. 3 model lacks pixel-locality. Second, we note significant differences in conditional-locality between Exp. 1 vs. 2 and 3. The length-generalizing Exp. 1 model exhibits strong conditional-locality at high noise, while the non-length-generalizing Exp. 2 and 3 models lack conditional-locality at high noise (scores near conditioned locations either fail to respond to any conditions or depend on several non-local conditions). At low noise, all models transition to pixel-local unconditional denoising, as in Kamb and Ganguli (2024); Niedoba et al. (2024). These experiments support our prediction that length generalization depends on pixel- and conditional-locality, and suggest that conditional-locality at high noise plays a particularly important role. Locality metrics are plotted in Figure 12, Figure 9 shows additional pixel locations, and Appendix F details the locality measurements. Figure 2 plots length generalization vs. conditioner locality for several models (different colors), each checkpointed early, mid, and late in training (different shapes). Details are in Section E.2.1. Length generalization and conditional locality are strongly correlated and can emerge together over the course of training.
@@
Experiment 2L (Figure 1) tests our theory via a direct causal intervention, wherein we design a local model architecture that explicitly enforces the local conditional scores defined in our theory. Specifically, we modify the architecture to process images as a grid of independent patches, each conditioned on the subset of locations within the patch. This design is conceptually related to local model architectures proposed in prior works like Li et al. (2023); Zheng et al. (2023); Cheng et al. (2023); Watters et al. (2019). We train the local model using the same conditioning as the failing Exp. 2 (labeling only a single object location), and find that the local architectural intervention causes it to length-generalize (Table 1). In fact, the local model trained on only one object can length-generalize up to 5 locations (Figure 8, Table 1). This supports the hypothesis that local conditional scores could be a causal mechanism for compositional generalization. Locality metrics and plots for Exp. 2L are shown in Figure 12 and 9. Additional details in Section E.3. We perform an analogous local intervention for Experiment 3 that similarly improves length-generalization as described in Section E.4.
@@
In Section E.5 (Figure 13 and Table 4), we explore length generalization in color-conditioned CLEVR, where we may expect compositional structure to exist in an appropriate feature space. We observe that length generalization is possible to some extent – e.g. a model trained on 1-5 objects can generalize to 7 colors.
SDXL investigation
Figure 4: SDXL pixel-space locality. (Left) An SDXL-generated image for the prompt “a beautiful photograph with a horse in the middle, a dog on the left, and a cat on the right,” with four analysis locations marked. (Center) Pixel gradient magnitude heatmaps at low, mid, and high noise for each location, showing pixel-locality (localized cross patterns) especially at low noise. (Right) Per-word conditional influence (ablation) over time at two pixel locations: the curves are nearly identical despite the pixels being in different spatial regions, indicating a lack of conditional-locality in pixel-space. Averaged over 10 seeds; additional seeds and prompts in Figure 14, spatial word-influence heatmaps in Figure 15.
Can local mechanisms help to explain compositional generalization in real-world diffusion text-to-image models? While length generalization served as a controlled and verifiable special-case of compositional generalization in our CLEVR experiments, we now return to the broader question of compositions of novel combinations of concepts. We begin to explore this question by studying compositional/local structure in both pixel- and feature-space in a pretrained (‘out-of-the-box’) SDXL model (Podell et al., 2023). In Figure 4, we investigate pixel-space locality using a prompt containing implicit location information (“middle”, “right”, “left”). We measure pixel-locality via pixel gradient magnitude (Jacobian heatmaps) and conditional-locality via per-word ablation influence, where each word’s token embedding is replaced with padding and the resulting change in the score is measured (averaged over 10 seeds; see Appendix G&F for details). We find that pixel-locality is present, particularly at low noise: the Jacobian heatmaps show localized cross-shaped patterns centered on each analysis pixel (reflecting the effective receptive field of the convolutional U-Net architecture (Luo et al., 2016)). However, conditional-locality is largely absent in pixel-space: the per-word influence profiles are nearly identical at different pixel locations (Figure 4, right), and spatial word-influence heatmaps are largely spatially uniform (Figures 14 and 15). Consistent with this lack of pixel-space conditional-locality, SDXL rarely succeeds in correctly composing all three animals at the specified locations (Figure 14). This lack of pixel-space conditional-locality motivates looking beyond pixel-space.
[…]
Figure 5: Feature-space disentanglement in SDXL. (Top left) Within- vs. between-category mean cosine similarity (F-LCS heuristic, Lemma 2) for hidden-state activations in the down, mid, and up blocks (at high noise) and for output spaces (latent/VAE, pixel). Feature-space representations are more disentangled (larger within/between gap) than output spaces. (Top right) Per-layer disentanglement ratio (within/between cosine similarity) across all transformer layers, showing an arch-shaped profile peaking near the U-net bottleneck. (Middle) $8\times 8$ cosine similarity heatmaps for 8 representative concepts across feature-space (down, mid, up blocks) and output spaces (latent, pixel). Block-diagonal category structure (animals, art styles, foods) is most pronounced in feature-space. (Bottom) Example SDXL compositions: concepts that are disentangled in feature-space compose more successfully. Methodology: 24 prompts across 4 categories, 10 seeds, hidden states (post-FFN activations from BasicTransformerBlock) at high noise; see Appendix G and Figure 16.
@@ 5 Discussion and future work
Whether local mechanisms can explain compositional generalization in real-world diffusion models remains an open question. The preliminary SDXL experiments in Section 4 are meant only to be suggestive, and much more work is needed to fully understand compositional generalization in modern text-to-image models. This setting presents several challenges. First, since we often don’t know what was in the training set, it is unclear which prompts are actually OOD (however, see Appendix H for a small exploratory study of a model trained on a known dataset). Second, some kinds of compositional structure, such as style+content, exist only in feature-space, requiring more complex studies of the model’s learned representation as in Figures 13 and 5, potentially relying on heuristics such as Lemma 2. Despite significant evidence for the existence of disentangled feature spaces (Chen et al., 2018; Kim and Mnih, 2018; Yang et al., 2023; Locatello et al., 2019; Zhang et al., 2023; Ilharco et al., 2022) and diffusion models’ ability to learn them in some cases Karras et al. (2019); Kotovenko et al. (2019); Gatys et al. (2016); Zhu et al. (2017), disentanglement in diffusion representations is still not fully understood. Future work could also exploit our finding of local conditional scores as a mechanism to improve compositional generalization. Our causal intervention in Experiment 2L shows that in a simple setting, enforcing an explicitly local architecture can improve generalization, suggesting that similar architectural, training, or inference-based interventions might be able to improve real-world models. Several existing methods can be viewed as implicit “local interventions”: for instance, layout-to-image methods that use explicit local constraints or biases (see Appendix A), or more generally, sparse attention architectures Child et al. (2019); Sun et al. (2022). Our theory helps to explain why these approaches are beneficial, and suggests more precise ways to target compositional generalization by specifically enforcing local conditional scores. When compositional structure exists only in feature-space, such interventions become more complex. The challenge becomes two-fold: we must first identify or attempt to induce feature-space transformations revealing the compositional structure, and then apply local interventions within the learned feature space: perhaps via sparsity-inducing regularization (such as L1) or explicitly-sparse architectures. Larger-scale studies on complex, real-world datasets are essential to clarify the challenges and explore opportunities to improve compositional generalization via local interventions.